Calculate the future value at the end of year 3 of an investment fund earning 12 percent annual interest and funded with the following end-of-year deposits: £750 in at the end of year 1, £850 and at the end of year 2 and £950 at the end of year 3.
1
Expert's answer
2014-07-03T08:44:49-0400
Future Value = Present Value * (1+r)^n

Present Value = C ÷ (1 + r)^n,
Where C - Cash flow at a period; r - rate of return and n - number of periods.

PV1= 750 ÷ 1.12 = 669.64
PV2 =850 ÷ 1.12^2 = 677.61
PV3= 950 ÷ 1.12^3 = 676.19
Total= 2023.44

FV =2023.44 * 1.12^3 = 2842.79